miguel's coin collection consists of dimes and quarters. if he has 53 coins worth $7.85, how many of each type of coin does he have?
Answer by Alan3354(69443)

miguel's coin collection consists of dimes and quarters. if he has 53 coins worth $7.85, how many of each type of coin does he have?
------------
Miguel must be very proud.
d + q = 53
10d+ 25q = 785
2d + 5q = 157
2d + 2q = 106
------------- Subtract
3q = 51
q = 17 quarters
d = 36 dimes
--------------
Keep it in a safe place